# Metastable Strings and Gravitational Waves in One-Scale Models

### v1

@@ 5 Results
Pulsar Timing Arrays (PTAs) are uniquely sensitive to gravitational waves (GWs) in the nanohertz frequency range. By using long observational baselines to monitor pulse arrival times from an array of millisecond pulsars, PTAs can detect a stochastic GW background (SGWB), which manifests as spatially correlated timing residuals across the sky. In 2023, the NANOGrav collaboration reported compelling evidence for such a signal, identifying quadrupolar spatial correlations consistent with the Hellings–Downs curve [Hellings:1983fr] and strongly supporting a GW origin [NANOGrav:2023gor, NANOGrav:2023icp, NANOGrav:2023hfp, NANOGrav:2023ctt, NANOGrav:2023hvm]. This landmark discovery was independently confirmed by EPTA [EPTA:2023fyk, EPTA:2023sfo, EPTA:2023akd, EPTA:2023gyr, EPTA:2023xxk, EuropeanPulsarTimingArray:2023egv], PPTA [Zic:2023gta, Reardon:2023zen, Reardon:2023gzh], and CPTA [Xu:2023wog]. While a population of inspiralling supermassive black hole binaries is the most anticipated astrophysical explanation, alternative cosmological sources remain viable. These include first-order phase transitions, cosmic superstrings [Ellis:2023tsl], and metastable cosmic strings [Buchmuller:2023aus, Antusch:2023zjk, Madge:2023dxc]. Notably, Bayesian analyses of the PTA data already disfavour standard, stable Nambu–Goto string networks, placing strong limits on the symmetry-breaking scale for such models. This motivates a deeper investigation into alternative string scenarios. Ongoing PTA measurements, alongside future probes such as space- and ground-based interferometers and atomic sensors, offer powerful tests of grand unified theories and early-universe dynamics [Buchmuller:2013lra, Dror:2019syi, Buchmuller:2019gfy, Chigusa:2020rks, Saad:2022mzu, Fu:2023mdu].
Our main result is summarised in Figure 3, which shows the regions of model parameter space that can consistently reproduce the PTA stochastic background [NANOGrav:2023gor, EPTA:2023fyk, Reardon:2023zen, Xu:2023wog]. If this stochastic gravitational wave background arises from the decay of cosmic strings, the signal’s amplitude and spectral shape constrain the rate of string decay per unit length, $\Gamma$. This rate is, in turn, related to the parameter $\kappa$ via the semiclassical decay formula [Monin:2008mp, Monin:2009ch, Leblond:2009fq],
@@
The NANOGrav 15-year analysis [NANOGrav:2023hvm] interprets the stochastic gravitational-wave signal in terms of two phenomenological models of metastable cosmic-string networks. The first, denoted meta-l, assumes that gravitational radiation is dominated by oscillating loops produced by the long-string network. The second, meta-ls, augments this picture by including a contribution from string segments: finite-length pieces of string bounded by monopole–antimonopole pairs. The principal physical distinction between these models lies in whether vibrational energy on the segments can efficiently radiate as gravitational waves. If unconfined gauge flux can escape from the segment endpoints, as in some microscopic realisations, then energy preferentially dissipates into gauge radiation and segment emission is suppressed, which realises the meta-l scenario. Conversely, if gauge flux is confined so that vibrational energy remains trapped on the segments, then segment emission can make an appreciable contribution and the behaviour corresponds to meta-ls.
The analysis of [NANOGrav:2023hvm] reports that both models can reproduce the observed signal, but shows a mild preference for the meta-ls fit over meta-l, as reflected in the reported Bayes factors. Both types of behaviour arise naturally within our electroweak-like metastable string framework. In the minimal, single-scale realisation, the dark $U(1)_{\text{IR}}$ remains unbroken at low energies and the magnetic flux associated with the string can escape from segment endpoints. Segment emission is therefore suppressed, which corresponds to meta-l. An equally straightforward variant occurs if the $U(1)_{\text{IR}}$ is Higgsed at a lower scale, in which case the flux is confined and segment emission mirrors the meta-ls prescription. The two NANOGrav model assumptions therefore correspond to two limiting cases within the same underlying microscopic setup.
@@
Figure 3: Regions of parameter space in the $(\beta,\sin^{2}\theta_{w})$ plane. Shaded in pink are points with $\kappa>8.5$; dark green corresponds to $8.5>\kappa>8$; light green to $8>\kappa>7.5$; and blue to $\kappa<7.5$. The grey contour shows the metastable region taken from Fig. 1 of Ref. [James:1992wb], based on an analysis of the spectrum of single particle fluctuations around the classical string background. The three panels correspond to different values of the $U(1)$ fine structure constant of $\alpha^{\prime}=g^{\prime 2}/4\pi$. An overlap between the metastable region and the $\kappa$-favoured band is visible for $\alpha^{\prime}\geq 3/4$, showing that the PTA-preferred range can be accommodated for moderately large $U(1)$ gauge couplings.
@@
In Figure 3, the regions shaded light and dark green yield values for $\kappa$ inside the range indicated in Eq. (5.2). The three panels of the figure show how these regions shift as a function of the parameter $\alpha^{\prime}$. Not all points in parameter space correspond to strings that are long-lived enough for the semiclassical decay treatment to apply. In particular, electroweak strings are only metastable in part of their parameter space. This is determined from the stability analysis performed in Ref. [James:1992wb], in which single particle fluctuations around the classical background given by the string profile functions are considered, and their spectrum is computed. If all fluctuation modes have positive energy, the string configuration is metastable and can decay semiclassically with a rate governed by $\kappa$. If, however, any fluctuation mode has negative energy, the configuration is unstable and decays promptly to particle excitations. The semiclassical description of the decay rate is not valid in this case.
The metastable region taken from the fluctuation analysis of Ref. [James:1992wb] is overlaid as a grey region in Figure 3. Only points lying inside this triangular region correspond to cosmic strings that are metastable, and can produce the gravitational wave signal under consideration. The triangular region is bounded from below by a dashed grey line. The string profile functions entering the stability analysis were only computed numerically for $\beta$ values above this bound, since this computation becomes more challenging for small $\beta$, when the two profile functions show significant variation over two different physical scales ($M_{\Phi}$ and $M_{Z}$) that are widely separated. It would be interesting to explore stability in the portion of parameter space below this line in future work.
@@
However, the $U(1)$ coupling $\alpha^{\prime}\sim 1$, and one could worry if the non-weakly coupled dynamics of the Abelian sector would invalidate the semiclassical approach based on expanding the path integral around the classical string solution. We do not expect that this is the case. To concentrate on the effects of $\alpha^{\prime}$ we can for simplicity decouple the non-Abelian sector. In the resulting Abelian Higgs model the string is an infinitely long-lived, localised field configuration that carries a conserved and divergenceless magnetic flux. Since no flux-free configuration can be continuously reached from the flux-carrying one, the string cannot decay into configurations where the flux is absent. Furthermore, as long as $\alpha_{\rm str}<\alpha_{\infty}=\pi$, which holds throughout the parameter range of interest, there are no alternative flux-carrying configurations with lower energy than the string. These conclusions rely solely on flux conservation and the energetics of Abelian vortex solutions, and do not depend on whether the Abelian Higgs model is weakly or strongly coupled. The string remains a valid configuration to expand around. Corrections in $\alpha^{\prime}$ will affect the form of the pre-factor in front of the exponent in the decay rate formula of Eq. (4.15) but we do not expect them to invalidate our conclusions. The pre-factor emerges from integrating out fluctuations around the bounce solution in the path integral and the relevant contributions are closed loops (bubble diagrams) involving propagators in the background of the bounce configurations. At one-loop level the prefactor was computed in [Monin:2008mp] which for our $SU(2)\times U(1)$ model results in,
@@ 6 Discussion
We work in a one-scale setting, where the defect physics is controlled by a single symmetry-breaking scale, the dark Higgs vacuum expectation value, $\eta$. This contrasts with multi-step breaking which can also lead to the formation of metastable strings $SU(2)\to U(1)\to\emptyset$, in which monopoles from the first transition can nucleate on strings formed at the second [Buchmuller:2019gfy, Buchmuller:2023aus, Shifman:2002yi, Chitose:2023dam, Chitose:2025cmt]. The two gauge symmetry breaking transitions occur at the scales $V$ and $v$ respectively. String tension is governed by the lower of the two scales, $\mu_{\rm str}\sim 2\pi v^{2}.$ These strings are metastable and can decay into monopole pairs by unwinding their field configurations via the $SU(2)$ gauge fields. Monopole configurations in this model are governed by the large vacuum expectation value, $M_{m}\sim 4\pi V/g.$ As shown in [Shifman:2002yi], the resulting string decay rate can be calculated semiclassically and the approach is robust in a thin-wall limit which relies on the hierarchy of the vacuum expectation values: $V\gg v$. To fit the semiclassical rate exponent $\kappa$ in these models into the range relevant for the PTA signal, one is forced however to forgo the large hierarchy of scales and consider the regime $V\to v$ where the two steps of the symmetry breaking pattern are not well-separated [Buchmuller:2023aus]. In this case, the thin-wall approximation is invalid and one has to resort to other means of finding the bounce solution and the bounce action. This program was carried out in Ref. [Chitose:2023dam] relying on the Shifman-Yung ansatz [Shifman:2002yi] to parameterise the tunnelling trajectory, however in the absence of the large hierarchy between the scales, $V\gg v$, there is no guarantee that the ansatz describes the dominant trajectory in field space [Shifman:2002yi]. It is also not apriori obvious that the corresponding string configuration is classically stable in the first place in the $V\sim v$ regime, and if not, it will decay classically and much faster than the required exponentially suppressed semiclassical decay. These features of two-scale models are different from the single-scale model setup considered in our paper where the string configuration itself is known to be classically stable, and its quantum mechanical decay rate can be reliably calculated in the thin-wall approach.

### v2

@@ 5 Results
Pulsar Timing Arrays (PTAs) are uniquely sensitive to gravitational waves (GWs) in the nanohertz frequency range. By using long observational baselines to monitor pulse arrival times from an array of millisecond pulsars, PTAs can detect a stochastic GW background (SGWB), which manifests as spatially correlated timing residuals across the sky. In 2023, the NANOGrav collaboration reported compelling evidence for such a signal, identifying quadrupolar spatial correlations consistent with the Hellings–Downs curve [40] and strongly supporting a GW origin [7, 47, 5, 6, 4]. This landmark discovery was independently confirmed by EPTA [10, 8, 9, 12, 11, 76], PPTA [87, 69, 68], and CPTA [84]. While a population of inspiralling supermassive black hole binaries is the most anticipated astrophysical explanation, alternative cosmological sources remain viable. These include first-order phase transitions, cosmic superstrings [32], and metastable cosmic strings [22, 13, 54]. Notably, Bayesian analyses of the PTA data already disfavour standard, stable Nambu–Goto string networks, placing strong limits on the symmetry-breaking scale for such models. This motivates a deeper investigation into alternative string scenarios. Ongoing PTA measurements, alongside future probes such as space- and ground-based interferometers and atomic sensors, offer powerful tests of grand unified theories and early-universe dynamics [19, 30, 20, 24, 71, 37].
Our main result is summarised in Figure 3, which shows the regions of model parameter space that can consistently reproduce the PTA stochastic background [7, 10, 69, 84]. If this stochastic gravitational wave background arises from the decay of cosmic strings, the signal’s amplitude and spectral shape constrain the rate of string decay per unit length, $\Gamma$. This rate is, in turn, related to the parameter $\kappa$ via the semiclassical decay formula [57, 58, 53],
@@
The NANOGrav 15-year analysis [4] interprets the stochastic gravitational-wave signal in terms of two phenomenological models of metastable cosmic-string networks. The first, denoted meta-l, assumes that gravitational radiation is dominated by oscillating loops produced by the long-string network. The second, meta-ls, augments this picture by including a contribution from string segments: finite-length pieces of string bounded by monopole–antimonopole pairs. The principal physical distinction between these models lies in whether vibrational energy on the segments can efficiently radiate as gravitational waves. If unconfined gauge flux can escape from the segment endpoints, as in some microscopic realisations, then energy preferentially dissipates into gauge radiation and segment emission is suppressed, which realises the meta-l scenario. Conversely, if gauge flux is confined so that vibrational energy remains trapped on the segments, then segment emission can make an appreciable contribution and the behaviour corresponds to meta-ls. It has shown that composite monopole-string structures in $SO(10)\to SU(5)\times U(1)$ breaking patterns produce configurations where monopoles connected by strings do not carry unconfined flux after electroweak breaking [55], providing a concrete realisation that favours the meta-ls scenario and yields improved fits to NANOGrav 15-year data.
The analysis of [4] reports that both models can reproduce the observed signal, but shows a mild preference for the meta-ls fit over meta-l, as reflected in the reported Bayes factors. Both types of behaviour arise naturally within our electroweak-like metastable string framework. In the minimal, single-scale realisation, the dark $U(1)_{\text{IR}}$ remains unbroken at low energies and the magnetic flux associated with the string can escape from segment endpoints. Segment emission is therefore suppressed, which corresponds to meta-l. An equally straightforward variant occurs if the $U(1)_{\text{IR}}$ is Higgsed at a lower scale, in which case the flux is confined and segment emission mirrors the meta-ls prescription. The two NANOGrav model assumptions therefore correspond to two limiting cases within the same underlying microscopic setup.
@@
Figure 3: Regions of parameter space in the $(\beta,\sin^{2}\theta_{w})$ plane. Shaded in pink are points with $\kappa>8.5$; dark green corresponds to $8.5>\kappa>8$; light green to $8>\kappa>7.5$; and blue to $\kappa<7.5$. The grey contour shows the metastable region taken from Fig. 1 of Ref. [46], based on an analysis of the spectrum of single particle fluctuations around the classical string background. The three panels correspond to different values of the $U(1)$ fine structure constant of $\alpha^{\prime}=g^{\prime 2}/4\pi$. An overlap between the metastable region and the $\kappa$-favoured band is visible for $\alpha^{\prime}\geq 3/4$, showing that the PTA-preferred range can be accommodated for moderately large $U(1)$ gauge couplings.
@@
In Figure 3, the regions shaded light and dark green yield values for $\kappa$ inside the range indicated in Eq. (5.2). The three panels of the figure show how these regions shift as a function of the parameter $\alpha^{\prime}$. Not all points in parameter space correspond to strings that are long-lived enough for the semiclassical decay treatment to apply. In particular, electroweak strings are only metastable in part of their parameter space. This is determined from the stability analysis performed in Ref. [46], in which single particle fluctuations around the classical background given by the string profile functions are considered, and their spectrum is computed. If all fluctuation modes have positive energy, the string configuration is metastable and can decay semiclassically with a rate governed by $\kappa$. If, however, any fluctuation mode has negative energy, the configuration is unstable and decays promptly to particle excitations. The semiclassical description of the decay rate is not valid in this case.
The metastable region taken from the fluctuation analysis of Ref. [46] is overlaid as a grey region in Figure 3. Only points lying inside this triangular region correspond to cosmic strings that are metastable, and can produce the gravitational wave signal under consideration. The triangular region is bounded from below by a dashed grey line. The string profile functions entering the stability analysis were only computed numerically for $\beta$ values above this bound, since this computation becomes more challenging for small $\beta$, when the two profile functions show significant variation over two different physical scales ($M_{\Phi}$ and $M_{Z}$) that are widely separated. It would be interesting to explore stability in the portion of parameter space below this line in future work.
@@
However, the $U(1)$ coupling $\alpha^{\prime}\sim 1$, and one could worry if the non-weakly coupled dynamics of the Abelian sector would invalidate the semiclassical approach based on expanding the path integral around the classical string solution. We do not expect that this is the case. To concentrate on the effects of $\alpha^{\prime}$ we can for simplicity decouple the non-Abelian sector. In the resulting Abelian Higgs model the string is an infinitely long-lived, localised field configuration that carries a conserved and divergenceless magnetic flux. Since no flux-free configuration can be continuously reached from the flux-carrying one, the string cannot decay into configurations where the flux is absent. Furthermore, as long as $\alpha_{\rm str}<\alpha_{\infty}=\pi$, which holds throughout the parameter range of interest, there are no alternative flux-carrying configurations with lower energy than the string. These conclusions rely solely on flux conservation and the energetics of Abelian vortex solutions, and do not depend on whether the Abelian Higgs model is weakly or strongly coupled. The string remains a valid configuration to expand around. Corrections in $\alpha^{\prime}$ will affect the form of the pre-factor in front of the exponent in the decay rate formula of Eq. (4.15) but we do not expect them to invalidate our conclusions. The pre-factor emerges from integrating out fluctuations around the bounce solution in the path integral and the relevant contributions are closed loops (bubble diagrams) involving propagators in the background of the bounce configurations. At one-loop level the prefactor was computed in [57] which for our $SU(2)\times U(1)$ model results in,
@@ 6 Discussion
One aspect of our general setup that we have not investigated in detail is the formation of the string network. A potential concern is whether the initial defect population is dominated by short string segments with monopole endpoints that collapse rapidly, rather than a long-lived network of “infinite” strings as we have assumed. However, at least in one-scale settings in the semilocal limit, it has been established that networks with significant populations of long strings do form, as reviewed in Refs. [3, 78, 2]. This supports our assumption that a metastable network persists until the semiclassical decay epoch in the parameter region we consider where the non-Abelian gauge coupling $g$ is small, without the need for additional inflationary stages to dilute monopoles. This economy contrasts with many multi-stage scenarios, where an intermediate inflationary period is invoked to dilute monopoles while preserving the long-string component.
We work in a one-scale setting, where the defect physics is controlled by a single symmetry-breaking scale, the dark Higgs vacuum expectation value, $\eta$. This contrasts with multi-step breaking which can also lead to the formation of metastable strings $SU(2)\to U(1)\to\emptyset$, in which monopoles from the first transition can nucleate on strings formed at the second [20, 22, 75, 25, 26]. The two gauge symmetry breaking transitions occur at the scales $V$ and $v$ respectively. String tension is governed by the lower of the two scales, $\mu_{\rm str}\sim 2\pi v^{2}.$ These strings are metastable and can decay into monopole pairs by unwinding their field configurations via the $SU(2)$ gauge fields. Monopole configurations in this model are governed by the large vacuum expectation value, $M_{m}\sim 4\pi V/g.$ As shown in [75], the resulting string decay rate can be calculated semiclassically and the approach is robust in a thin-wall limit which relies on the hierarchy of the vacuum expectation values: $V\gg v$. To fit the semiclassical rate exponent $\kappa$ in these models into the range relevant for the PTA signal, one is forced however to forgo the large hierarchy of scales and consider the regime $V\to v$ where the two steps of the symmetry breaking pattern are not well-separated [22]. In this case, it is not a priori clear that the corresponding string configuration is classically stable in the regime $V\sim v$; if it is not, the string would undergo rapid classical decay, overwhelming the exponentially suppressed semiclassical decay process. These features of two-scale models are different from the single-scale model setup considered in our paper where the string configuration itself is known to be classically stable, and its quantum mechanical decay rate can be reliably calculated.